OpenStudy (anonymous):

7/8 divided by 11/12=

OpenStudy (anonymous):

\[\frac{ 7 }{ 8 }\div \frac{ 11 }{ 12 }=\]

OpenStudy (anonymous):

\[\frac{ 7 }{ 8 }\div \frac{ 11 }{ 12 }=\frac{ 7 }{ 8 }\times \frac{ 12 }{ 11 }=\]

OpenStudy (anonymous):